# CHEMOUT: CHEMical complexity in star-forming regions of the OUTer Galaxy

### v2

@@ 1 Introduction
The outer Galaxy (OG), namely the portion of the Galactic disk beyond the solar circle, is located at galactocentric distances ($R_{\text{GC}}$) approximately between 9 and 24 kpc, and it shows chemical properties significantly different from those of the inner Galaxy. Several chemical galactocentric gradient studies have been conducted, mapping the differences in elemental composition between the inner and outer Milky Way (e.g., C, N, O, Fe), using various tracers (i.e., H II regions, classical Cepheids, young OB stars, and planetary nebulae, among others). These works, focused on observational trends (e.g., luck2018cepheid; bragancca2019radial; mendez2022gradients) and chemodynamical studies (jacobson2016gaia; stanghellini2018galactic; palla2020chemical; santos2021ambre), indicate that the overall metallicity in the OG is lower than the solar one (e.g., a factor of four lower at $R_{\text{GC}}=19\penalty 10000\$kpc; shimonishi2021detection). The elemental abundances of carbon, oxygen, and nitrogen - the three most abundant elements in the Universe after hydrogen and helium - decrease as a function of $R_{\text{GC}}$ (e.g., esteban2017radial; arellano2020galactic), similarly to other heavier elements, following the radial metallicity gradient. This observational result is linked to the traditional view of the OG as being poorer in molecules than the solar neighborhoods and the inner Galaxy. Moreover, based on the aforementioned scarcity of elements crucial for the formation of complex molecules and (rocky) planets, chemical evolution models excluded most of the OG from the so-called Galactic habitable zone (GHZ); that is, the region in the Milky Way where the probability for the development of habitable planets is highest: the optimal region is predicted to be an annulus of the Galactic disk from $\sim 7\penalty 10000\$kpc extended up to $R_{\text{GC}}\sim 9\penalty 10000\$kpc, with maximum habitability at $\sim 8$ kpc (spitoni2014galactic; spitoni2017evolution). This led to the consideration of the OG not being suitable for the formation of planetary systems in which Earth-like planets could be born and might be capable of sustaining life (prantzos2008galactic; ramirez2010possible).
@@
Recent observational results challenged this prediction. First, the occurrence of Earth-like planets does not seem to depend on metallicity (e.g., maliuk2020spatial; mulders2018exoplanet). Second, the low rate of disruptive events in the OG suggests that it might be more favorable than the inner Galaxy for preserving life on potentially habitable planets (e.g., piran2014possible; vukotic2016grandeur). Third, the detection of several complex organic molecules (COMs; e.g., CH3OH, CH3CCH, CH3OCHO, CH3OCH3) of prebiotic interest in low-metallicity environments, such as star-forming regions in the OG (shimonishi2021detection; bernal2021methanol; fontani2022chemout; fontani2024chemout; koelemay2025no) and the Magellanic Clouds (sewilo2018detection; sewilo2022alma; shimonishi2023detection; golshan2024high). Overall, these results suggest that environments prone to the presence of molecular precursors of life and Earth-like planets might be widespread in the Milky Way, including the OG.
@@
The project CHEMical complexity in the outer Galaxy (CHEMOUT; fontani2022chemout fontani2022chemout, fontani2022chemoutII, colzi2022chemout colzi2022chemout, fontani2024chemout fontani2024chemout), performed with the IRAM 30 m single-dish telescope, is aimed at deriving the molecular composition in 35 star-forming regions of the OG ($R_{\text{GC}}\sim 9-24\penalty 10000\$kpc). The main results obtained in the previous works of the CHEMOUT project are (1) both the molecular detection rates and the molecular abundances of some organic species –namely HCO, H2CO, and CH3OH– do not vary significantly with $R_{\text{GC}}$ (fontani2022chemout fontani2022chemout, fontani2022chemoutII); (2) the 14N/15N ratio as a function of $R_{\text{GC}}$ is consistent with chemical-evolution models in which 15N comes from novae (colzi2022chemout); (3) the high-angular-resolution maps of the target with the highest $R_{\text{GC}}$ (WB89-670 at $R_{\text{GC}}\sim 23.4$ kpc) reveal chemical differentiation at core scales and, above all, a [C/H] abundance higher than expected at such a high galactocentric distance (fontani2024chemout).
@@
The results obtained so far indicate a lack of decrease in the inspected molecular abundances, which, however, was limited to three species (HCO, H2CO, and CH3OH) and only 15 sources. In this paper, we analyze the molecules detected in more than $60\%$ of the CHEMOUT sample according to fontani2022chemout, namely HCN, HCO+, c-C3H2, H13CO+, HCO, and SO, and revise the molecular abundances of H2CO and CH3OH from fontani2022chemoutII, using H2 column densities from the new NIKA2 dust continuum maps. In Sect. 2, we describe the observational data used in this paper. In Sect. 3, we describe the methods implemented in the data analysis. The results are presented and discussed in Sect. 4 and 5. The conclusions and the outlooks are summarized in Sect. 6.
@@ 2.1 IRAM 30 m observations
This work is based on the observations done for the CHEMOUT project, described in Papers I and II (fontani2022chemout fontani2022chemout, fontani2022chemoutII), performed with the Institut de RadioAstronomie Millimétrique (IRAM) 30 m telescope. The sample of sources analyzed in this paper is composed of the 35 star-forming regions, presented in Table 1 of fontani2022chemout. In some spectra, two velocity features are identified, and the sources listed in Table 5 in this paper include both velocity components. In the analysis conducted, we used spectral windows at 3 and 2 mm, collected with the EMIR receiver and described in Table 1. The spectral windows of the project 004-18 (see Table 1) have been optimized to observe, at 3 mm, the J=10J=1-0 transitions of 13C and 15N isotopologues of HCN and HNC and, at 2 mm, the J=21J=2-1 line of DNC. The spectral windows observed in the project 042-21, at 3 and 2 mm, have been designed to observe several methanol and formaldehyde lines. For further details (e.g., pointing and focus, and weather conditions), we refer the reader to fontani2022chemout (fontani2022chemout, fontani2022chemoutII).
@@ 2.2 NIKA2 bolometer observations
Millimeter continuum emission of 31 out of the 35 CHEMOUT targets was observed with the New IRAM KID Arrays 2 (NIKA2; perotto2020) bolometer of the IRAM 30 m telescope in Pico Veleta, Spain. Observations were taken during the winter 2021 and 2023 pool seasons (February). The observations cover an almost circular region with a radius of $\sim 9$ arcmin, centered at the positions where the IRAM 30 m spectra have been observed (Table 5). The dual-band capability of NIKA2 allowed us to simultaneously observe the dust’s thermal continuum emission at 2.0 and 1.2 mm (150 GHz and 260 GHz, respectively). The sensitivity achieved was $\sim 1.5$ mJy/beam at 1.2 mm and 0.5 mJy/beam at 2.0 mm. A more complete description of the observations and a detailed analysis of the maps will be presented in a forthcoming paper of the CHEMOUT series (Fontani et al. in prep). Here, we have used the maps only to derive the H2 column densities (Sect. 3.2) to compute the molecular fractional abundances with respect to H2 (Sect. 4.2).
@@ 3.1 Line fitting
The spectra were fit using the MAdrid Data CUBe Analysis (MADCUBA [^2] MADCUBA is a software developed at the Madrid Center of Astrobiology (CAB, CSIC-INTA) that enables the visualization and analysis of single spectra and data cubes: https://cab.inta-csic.es/madcuba/; martin2019spectral) software. The transitions were identified using the Spectral Line Identification and LTE Modelling (SLIM) tool of MADCUBA, which makes use of the Jet Propulsion Laboratory (JPL; pickett1998submillimeter) and Cologne Database for Molecular Spectroscopy (CDMS; muller2001cologne) catalogs. The lines were fit using the AUTOFIT function of SLIM. This function produces the Gaussian synthetic spectrum that best matches the observed spectrum, assuming local thermodynamic equilibrium (LTE) conditions and the same fitting input parameters for all transitions of a given molecular species, which are total column density, $N_{\text{tot}}$; excitation temperature, $T_{\text{ex}}$; radial systemic velocity of the source, $V_{\text{LSR}}$; full width at half-maximum of the line, FWHM; and angular size of the emission, $\theta_{\text{S}}$.
@@
For the source size, we assumed that the molecular emission fills the telescope beam, except for methanol and formaldehyde, where an estimation of the emission size was made in fontani2022chemoutII and updated in this paper as discussed in Sect. 4.1.1. Such an assumption is justified by the spatial extent of the emission of the dust (see Figs. 4, 5, 6), as well as by the updated H2CO and CH3OH emission sizes shown in Table 8. The excitation temperature, whenever an estimate was not possible through the fit, has been fixed to the one shown in Table 5, which was estimated from the methanol lines for the 15 targets observed in fontani2022chemoutII. For the remaining sources, we adopted an average $T_{\text{ex}}$ of 10.7 K, derived from those sources for which methanol transitions were available. This approximation was necessary because it was not possible to estimate the temperature individually for each source based on the observed transitions. To evaluate the impact of this assumption, we computed $N_{\text{tot}}$ in the range of excitation temperatures derived from methanol ($7\leq T_{\text{ex}}\leq 15\penalty 10000\ K$; see Table 5). In this $T_{\text{ex}}$ range, the densities vary from the values shown in Tables 7 and 8 by a factor of two at most for the species studied (lowest for HCN, HCO+, and H13CO+; highest for SO).
@@
As a detection criterion, we considered a line as detected if the signal-to-noise ratio (S/N) is greater than three (an estimation of the rms noise is given in fontani2022chemout), while the detection is considered as tentative for $2.5\leq\text{S/N}\leq 3$. If a line is undetected (e.g., S/N $\leq 2.5$), MADCUBA provides a $3\sigma$ upper limit estimate to the line’s integrated intensity from Eq. (29) of martin2019spectral. The upper limit on the integrated intensity is converted into a column-density upper limit using the SLIM parameters $T_{\text{ex}}$ and $\theta_{\text{S}}$. For these estimates, the average FWHM for each species across the sample was adopted (Sect. 5.2). This approach assumes that different molecules trace gases with different physical, kinematic, and excitation conditions.
@@ 3.2 Continuum analysis
The column densities of H2, $N_{\text{H}_{2}}$ were derived from the dust-continuum-emission maps. In the scales (and densities) traced by our observations, the thermal dust emission is expected to be optically thin. Under this assumption, we can use the following equation to derive $N_{\text{H}_{2}}$ from the continuum flux density (e.g., battersby2011characterizing):
@@ 4.1 Detection summary
A comprehensive molecular identification was performed over the whole sample to identify the species detected in the greatest number of targets. This allowed us to analyze abundance gradients as a function of galactocentric distance with the highest possible statistics. Over 30 molecular species have been detected, and the most frequently detected ones are HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). The detection statistics over these species are already listed in fontani2022chemout. We also included H2CO (18/18) and CH3OH (17/18) among the species considered. The analysis of the HCO, H2CO, and CH3OH gradients was conducted by fontani2022chemoutII; however, here, we present updated trends based both on the revised fits performed with MADCUBA (Sect. 4.1.1) and on our new H2 column-density estimates from NIKA2 observations. The spectroscopic parameters of the transitions detected in these molecules are displayed in Table 3. In Appendix C, the best-fit parameters obtained from the analysis with MADCUBA, and the spectra of the fit transitions for HCN, H13CO+, and SO are shown in detail. The spectra of c-C3H2 and HCO+ are shown in fontani2022chemout, and the ones of HCO, H2CO, and CH3OH in fontani2022chemoutII. The results regarding the molecular column densities (e.g., their galactocentric gradients) are provided in Appendix D.
@@
In Table 5, the $N_{\text{H}_{2}}$ values calculated assuming a gas-to-dust ratio from giannetti2017galactocentric are presented. These values are used hereafter for the analysis. All the data, results, and discussions concerning fractional abundances with respect to H2 are based on these estimates, rather than on those obtained assuming a fixed $\gamma=100$. The NIKA2 dust continuum maps and the galactocentric gradients for H2 column densities are shown in Appendix B.
@@
Six sources (19383+2711, 19571+3113, WB89-006, WB89-060, WB89-080, and WB89-380) show multiple velocity features also reported in fontani2022chemout. Four of them (19383+2711, 19571+3113, WB89-006, and WB89-380) exhibit double-peak emission in most of the observed lines. Because the peak velocities of these two features are consistent in the different transitions in which they are detected, they probably arise from two clumps within the telescope beam that have an angular separation smaller than the beam itself. This is confirmed by the NIKA2 maps (Figs. 4 - 6), for 19383+2711 and 19571+3113, showing two different emission components in the beam of the IRAM 30 m telescope. WB89-006 and WB89-380 do not show a clear double peak in the continuum maps, but the two clumps could be non-resolved with NIKA2 or aligned along the line of sight. In principle, each of these clumps may have distinct physical and chemical properties and will therefore be analyzed separately in the following sections. For these, the peak velocities were estimated using the $V_{\text{LSR}}$ of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, and we assumed the same Galactic longitude, $\ell$, for both clumps, to estimate their galactocentric radius, $R_{\text{GC,}}$ and heliocentric distance, $d$. The remaining two sources (WB89-060 and WB89-080) are characterized by a double peak, but only in some more intense and potentially optically thick lines (e.g., HCN and HCO+) and do not show coherence in the peak velocities. Consequently, in these cases, the double-peaked line shapes are more likely due to high-optical-depth effects, such as (self-)absorption, rather than multiple velocity features. These have thus been analyzed as single sources, and the absorbed lines discarded from the analysis. Moreover, some transitions showed non-Gaussian, high-velocity wings. We decided to limit our analysis to the narrower component of the molecular emissions related to the line emission around the intensity peak, and thus arising from the bulk emission of the cores.

### v3

# arXiv:2509.26556v3[astro-ph.GA] 02 Mar 2026

@@ 1 Introduction
The outer Galaxy (OG), namely the portion of the Galactic disk beyond the solar circle, is located at galactocentric distances ($R_{\text{GC}}$) approximately between 9 and 24 kpc, and it shows chemical properties significantly different from those of the inner Galaxy. Several chemical galactocentric gradient studies have been conducted, mapping the differences in elemental composition between the inner and outer Milky Way (e.g., C, N, O, Fe), using various tracers (i.e., H II regions, classical Cepheids, young OB stars, and planetary nebulae, among others). These works, focused on observational trends (e.g., Luck 2018; Bragança et al. 2019; Méndez-Delgado et al. 2022) and chemodynamical studies (Jacobson et al. 2016; Stanghellini and Haywood 2018; Palla et al. 2020; Santos-Peral et al. 2021), indicate that the overall metallicity in the OG is lower than the solar one (e.g., a factor of four lower at $R_{\text{GC}}=19\penalty 10000\$kpc; Shimonishi et al. 2021). The elemental abundances of carbon, oxygen, and nitrogen - the three most abundant elements in the Universe after hydrogen and helium - decrease as a function of $R_{\text{GC}}$ (e.g., Esteban et al. 2017; Arellano-Córdova et al. 2020), similarly to other heavier elements, following the radial metallicity gradient. This observational result is linked to the traditional view of the OG as being poorer in molecules than the solar neighborhoods and the inner Galaxy. Moreover, based on the aforementioned scarcity of elements crucial for the formation of complex molecules and (rocky) planets, chemical evolution models excluded most of the OG from the so-called Galactic habitable zone (GHZ); that is, the region in the Milky Way where the probability for the development of habitable planets is highest: the optimal region is predicted to be an annulus of the Galactic disk from $\sim 7\penalty 10000\$kpc extended up to $R_{\text{GC}}\sim 9\penalty 10000\$kpc, with maximum habitability at $\sim 8$ kpc (Spitoni et al. 2014, 2017). This led to the consideration of the OG not being suitable for the formation of planetary systems in which Earth-like planets could be born and might be capable of sustaining life (Prantzos 2008; Ramírez et al. 2010).
@@
Recent observational results challenged this prediction. First, the occurrence of Earth-like planets does not seem to depend on metallicity (e.g., Maliuk and Budaj 2020; Mulders et al. 2018). Second, the low rate of disruptive events in the OG suggests that it might be more favorable than the inner Galaxy for preserving life on potentially habitable planets (e.g., Piran and Jimenez 2014; Vukotić et al. 2016). Third, the detection of several complex organic molecules (COMs; e.g., CH3OH, CH3CCH, CH3OCHO, CH3OCH3) of prebiotic interest in low-metallicity environments, such as star-forming regions in the OG (Shimonishi et al. 2021; Bernal et al. 2021; Fontani et al. 2022a, 2024; Koelemay and Ziurys 2025) and the Magellanic Clouds (Sewiło et al. 2018, 2022; Shimonishi et al. 2023; Golshan et al. 2024). Overall, these results suggest that environments prone to the presence of molecular precursors of life and Earth-like planets might be widespread in the Milky Way, including the OG.
@@
The project CHEMical complexity in the outer Galaxy (CHEMOUT; Fontani et al. 2022a, 2022b, Colzi et al. 2022, Fontani et al. 2024), performed with the IRAM 30 m single-dish telescope, is aimed at deriving the molecular composition in 35 star-forming regions of the OG ($R_{\text{GC}}\sim 9-24\penalty 10000\$kpc). The main results obtained in the previous works of the CHEMOUT project are (1) both the molecular detection rates and the molecular abundances of some organic species –namely HCO, H2CO, and CH3OH– do not vary significantly with $R_{\text{GC}}$ (Fontani et al. 2022a, 2022b); (2) the 14N/15N ratio as a function of $R_{\text{GC}}$ is consistent with chemical-evolution models in which 15N comes from novae (Colzi et al. 2022); (3) the high-angular-resolution maps of the target with the highest $R_{\text{GC}}$ (WB89-670 at $R_{\text{GC}}\sim 23.4$ kpc) reveal chemical differentiation at core scales and, above all, a [C/H] abundance higher than expected at such a high galactocentric distance (Fontani et al. 2024).
@@
The results obtained so far indicate a lack of decrease in the inspected molecular abundances, which, however, was limited to three species (HCO, H2CO, and CH3OH) and only 15 sources. In this paper, we analyze the molecules detected in more than $60\%$ of the CHEMOUT sample according to Fontani et al. (2022a), namely HCN, HCO+, c-C3H2, H13CO+, HCO, and SO, and revise the molecular abundances of H2CO and CH3OH from Fontani et al. (2022b), using H2 column densities from the new NIKA2 dust continuum maps. In Sect. 2, we describe the observational data used in this paper. In Sect. 3, we describe the methods implemented in the data analysis. The results are presented and discussed in Sect. 4 and 5. The conclusions and the outlooks are summarized in Sect. 6.
@@ 2.1 IRAM 30 m observations
This work is based on the observations done for the CHEMOUT project, described in Papers I and II (Fontani et al. 2022a, 2022b), performed with the Institut de RadioAstronomie Millimétrique (IRAM) 30 m telescope. The sample of sources analyzed in this paper is composed of the 35 star-forming regions, presented in Table 1 of Fontani et al. (2022a). In some spectra, two velocity features are identified, and the sources listed in Table 5 in this paper include both velocity components. In the analysis conducted, we used spectral windows at 3 and 2 mm, collected with the EMIR receiver and described in Table 1. The spectral windows of the project 004-18 (see Table 1) have been optimized to observe, at 3 mm, the J=10J=1-0 transitions of 13C and 15N isotopologues of HCN and HNC and, at 2 mm, the J=21J=2-1 line of DNC. The spectral windows observed in the project 042-21, at 3 and 2 mm, have been designed to observe several methanol and formaldehyde lines. For further details (e.g., pointing and focus, and weather conditions), we refer the reader to Fontani et al. (2022a, 2022b).
@@ 2.2 NIKA2 bolometer observations
Millimeter continuum emission of 31 out of the 35 CHEMOUT targets was observed with the New IRAM KID Arrays 2 (NIKA2; Perotto et al. 2020) bolometer of the IRAM 30 m telescope in Pico Veleta, Spain. Observations were taken during the winter 2021 and 2023 pool seasons (February). The observations cover an almost circular region with a radius of $\sim 9$ arcmin, centered at the positions where the IRAM 30 m spectra have been observed (Table 5). The dual-band capability of NIKA2 allowed us to simultaneously observe the dust’s thermal continuum emission at 2.0 and 1.2 mm (150 GHz and 260 GHz, respectively). The sensitivity achieved was $\sim 1.5$ mJy/beam at 1.2 mm and 0.5 mJy/beam at 2.0 mm. A more complete description of the observations and a detailed analysis of the maps will be presented in a forthcoming paper of the CHEMOUT series (Fontani et al. in prep). Here, we have used the maps only to derive the H2 column densities (Sect. 3.2) to compute the molecular fractional abundances with respect to H2 (Sect. 4.2).
@@ 3.1 Line fitting
The spectra were fit using the MAdrid Data CUBe Analysis (MADCUBA [^2] MADCUBA is a software developed at the Madrid Center of Astrobiology (CAB, CSIC-INTA) that enables the visualization and analysis of single spectra and data cubes: https://cab.inta-csic.es/madcuba/; Martín et al. 2019) software. The transitions were identified using the Spectral Line Identification and LTE Modelling (SLIM) tool of MADCUBA, which makes use of the Jet Propulsion Laboratory (JPL; Pickett et al. 1998) and Cologne Database for Molecular Spectroscopy (CDMS; Müller et al. 2001) catalogs. The lines were fit using the AUTOFIT function of SLIM. This function produces the Gaussian synthetic spectrum that best matches the observed spectrum, assuming local thermodynamic equilibrium (LTE) conditions and the same fitting input parameters for all transitions of a given molecular species, which are total column density, $N_{\text{tot}}$; excitation temperature, $T_{\text{ex}}$; radial systemic velocity of the source, $V_{\text{LSR}}$; full width at half-maximum of the line, FWHM; and angular size of the emission, $\theta_{\text{S}}$.
@@
For the source size, we assumed that the molecular emission fills the telescope beam, except for methanol and formaldehyde, where an estimation of the emission size was made in Fontani et al. (2022b) and updated in this paper as discussed in Sect. 4.1.1. Such an assumption is justified by the spatial extent of the emission of the dust (see Figs. 4, 5, 6), as well as by the updated H2CO and CH3OH emission sizes shown in Table 8. The excitation temperature, whenever an estimate was not possible through the fit, has been fixed to the one shown in Table 5, which was estimated from the methanol lines for the 15 targets observed in Fontani et al. (2022b). For the remaining sources, we adopted an average $T_{\text{ex}}$ of 10.7 K, derived from those sources for which methanol transitions were available. This approximation was necessary because it was not possible to estimate the temperature individually for each source based on the observed transitions. To evaluate the impact of this assumption, we computed $N_{\text{tot}}$ in the range of excitation temperatures derived from methanol ($7\leq T_{\text{ex}}\leq 15\penalty 10000\ K$; see Table 5). In this $T_{\text{ex}}$ range, the densities vary from the values shown in Tables 7 and 8 by a factor of two at most for the species studied (lowest for HCN, HCO+, and H13CO+; highest for SO).
@@
As a detection criterion, we considered a line as detected if the signal-to-noise ratio (S/N) is greater than three (an estimation of the rms noise is given in Fontani et al. 2022a), while the detection is considered as tentative for $2.5\leq\text{S/N}\leq 3$. If a line is undetected (e.g., S/N $\leq 2.5$), MADCUBA provides a $3\sigma$ upper limit estimate to the line’s integrated intensity from Eq. (29) of Martín et al. (2019). The upper limit on the integrated intensity is converted into a column-density upper limit using the SLIM parameters $T_{\text{ex}}$ and $\theta_{\text{S}}$. For these estimates, the average FWHM for each species across the sample was adopted (Sect. 5.2). This approach assumes that different molecules trace gases with different physical, kinematic, and excitation conditions.
@@ 3.2 Continuum analysis
The column densities of H2, $N_{\text{H}_{2}}$ were derived from the dust-continuum-emission maps. In the scales (and densities) traced by our observations, the thermal dust emission is expected to be optically thin. Under this assumption, we can use the following equation to derive $N_{\text{H}_{2}}$ from the continuum flux density (e.g., Battersby et al. 2011):
@@ 4.1 Detection summary
A comprehensive molecular identification was performed over the whole sample to identify the species detected in the greatest number of targets. This allowed us to analyze abundance gradients as a function of galactocentric distance with the highest possible statistics. Over 30 molecular species have been detected, and the most frequently detected ones are HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). The detection statistics over these species are already listed in Fontani et al. (2022a). We also included H2CO (18/18) and CH3OH (17/18) among the species considered. The analysis of the HCO, H2CO, and CH3OH gradients was conducted by Fontani et al. (2022b); however, here, we present updated trends based both on the revised fits performed with MADCUBA (Sect. 4.1.1) and on our new H2 column-density estimates from NIKA2 observations. The spectroscopic parameters of the transitions detected in these molecules are displayed in Table 3. In Appendix C, the best-fit parameters obtained from the analysis with MADCUBA, and the spectra of the fit transitions for HCN, H13CO+, and SO are shown in detail. The spectra of c-C3H2 and HCO+ are shown in Fontani et al. (2022a), and the ones of HCO, H2CO, and CH3OH in Fontani et al. (2022b). The results regarding the molecular column densities (e.g., their galactocentric gradients) are provided in Appendix D.
@@
In Table 5, the $N_{\text{H}_{2}}$ values calculated assuming a gas-to-dust ratio from Giannetti et al. (2017) are presented. These values are used hereafter for the analysis. All the data, results, and discussions concerning fractional abundances with respect to H2 are based on these estimates, rather than on those obtained assuming a fixed $\gamma=100$. The NIKA2 dust continuum maps and the galactocentric gradients for H2 column densities are shown in Appendix B.
@@
Six sources (19383+2711, 19571+3113, WB89-006, WB89-060, WB89-080, and WB89-380) show multiple velocity features also reported in Fontani et al. (2022a). Four of them (19383+2711, 19571+3113, WB89-006, and WB89-380) exhibit double-peak emission in most of the observed lines. Because the peak velocities of these two features are consistent in the different transitions in which they are detected, they probably arise from two clumps within the telescope beam that have an angular separation smaller than the beam itself. This is confirmed by the NIKA2 maps (Figs. 4 - 6), for 19383+2711 and 19571+3113, showing two different emission components in the beam of the IRAM 30 m telescope. WB89-006 and WB89-380 do not show a clear double peak in the continuum maps, but the two clumps could be non-resolved with NIKA2 or aligned along the line of sight. In principle, each of these clumps may have distinct physical and chemical properties and will therefore be analyzed separately in the following sections. For these, the peak velocities were estimated using the $V_{\text{LSR}}$ of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, and we assumed the same Galactic longitude, $\ell$, for both clumps, to estimate their galactocentric radius, $R_{\text{GC,}}$ and heliocentric distance, $d$. The remaining two sources (WB89-060 and WB89-080) are characterized by a double peak, but only in some more intense and potentially optically thick lines (e.g., HCN and HCO+) and do not show coherence in the peak velocities. Consequently, in these cases, the double-peaked line shapes are more likely due to high-optical-depth effects, such as (self-)absorption, rather than multiple velocity features. These have thus been analyzed as single sources, and the absorbed lines discarded from the analysis. Moreover, some transitions showed non-Gaussian, high-velocity wings. We decided to limit our analysis to the narrower component of the molecular emissions related to the line emission around the intensity peak, and thus arising from the bulk emission of the cores.